# One Stone, Four Birds: A Comprehensive Solution for QA System Using Supervised Contrastive Learning

### v1

@@ I Introduction
Figure 1: Our proposed workflow to achieve a comprehensive and adaptive QA system. First, an initial model is trained using current known data (Pre). Then, during daily usage, an inquiry will first be judged to be an answerable input or not (T-2): if so, proceed to the intent detection (T-1) and retrieve the answer; if not, the unknown inquiry will be temporally saved and be clustered later (T-3). Finally, the initial model is re-trained with newly discovered data to obtain the classification ability to new intents (T-4).
@@ 1. $\bullet$
    Externally, or at the system level: Current QA systems typically adopt a “training-as-complete” policy, meaning that once the model has finished training, its knowledge is not updated. This approach leaves the system vulnerable to unseen, unknown inputs (as out-of-domain, OOD inputs [11]). To maintain the robustness and security of the model, there should be mechanisms to handle these inputs. Furthermore, even when new data is collected to upgrade the system, the common practice is to re-train a new model from scratch, leading to a waste of labor and resources.
@@ 4. T-4
Figure 2: Diagrams of SCL’s representation optimization procedure, and solutions for each task. Left: the training and optimization effect by SCL, which gathers same-intent text features together and pushes different features further apart, providing an underlying embedding space for following tasks. Right: the classification/detection processes of T-1 to T-4 after SCL. (Together with Figure 1 for better understanding)
More specifically, the advantage of SCL over CE fine-tuning lies in its ability to enhance both learning efficiency and feature separation. As shown in Figure 2-b), SCL employs a simple but clear optimization target: pulling same-label text features together and pushing different ones further apart, resulting in an intra-class compact and inter-class scattered, well-organized embedding space. This clear separation between clusters facilitates easier in-domain (IND) intent classification and IND/OOD separation and detection (Figure2, T-1/2). Additionally, this structure also supports the aggregation of currently unknown but semantically similar features, creating a conducive environment for new intent discovery and continual learning (Figure 2, T-3/4). Consequently, even with the simplest algorithms in downstream tuning, all task objectives can be achieved effortlessly and efficiently.
@@
    Our experiments demonstrate that, despite its conciseness and simplicity, the proposed solution significantly improves global feature optimization efficiency and outperforms state-of-the-art methods across all tasks, complementing the deficiencies of current QA systems.
@@ II Related Works
II-A Representation Learning in Intent Classification Model Building
Early history of representation learning
Representation learning (RepL), or feature learning, is a crucial process in building machine learning models. Good RepL methods capture the intrinsic patterns and structures of the data, thus providing a solid foundation for downstream tasks.
Initially, RepL research focused on manually extracting meaningful features from raw data, with preliminary methods including TF-IDF [4] and BoW [5]. Subsequently, the development of word embeddings extracted via neural networks automated the feature extraction process [6, 7]. More recently, with the advent of large-scale encoders like BERT [8], pre-trained on extensive corpora, obtaining high-quality text embeddings has become much easier. The most common approach is to use BERT as the base encoder, with CE fine-tuning as the RepL method to refine the text embedding distribution.
Drawbacks of BERT+CE mode
The above BERT+CE approach is also widely adopted in QA systems [16, 10, 1, 17]. However, this method can be suboptimal for feature optimization from two main reasons. Firstly, from a global perspective, [18, 12] reported that the overall text embedding distribution generated by BERT is anisotropic. This means that most embeddings are spatially distributed in an elongated cone-like space rather than uniformly in a sphere, leading to high similarity between any two sentence representations and posing challenges for downstream tasks. Secondly, from a local perspective, CE fine-tuning results in a “loose” feature arrangement [13, 14], where the boundaries between different intent clusters are barely distinguishable. This not only reduces the performance of known, IND classification, but also makes the model prone to confusing IND inputs with unseen but similar OOD inputs.
@@ II-A Representation Learning in Intent Classification Model Building
As discussed above, RepL research has gradually shifted focus from feature extraction to feature optimization, with one of the most popular method series being the contrastive learning (ConL). The earliest ConL can be traced back to supervised metric learning (MeL) [19] used in image classification tasks, which established the fundamental strategy of “gathering similar examples together and pushing dissimilar ones further apart”. This strategy helps generate compact feature clusters efficiently. Subsequently, unsupervised ConL (UConL) [20, 21] was proposed, sharing the same strategy but aiming to improve the overall representation uniformity. Finally, SCL [14] incorporates the advantages of both MeL and UConL, thus having the potential to address both global and local feature optimization issues.
@@
However, ConL has been mainly examined in computer vision (CV) field. In NLP field, there have been limited applications, primarily in natural language inference (NLI) tasks using UConL methods [22, 23]. A boarder exploration of SCL’s learning effect across various tasks, including its application in QA systems, remains unaddressed.
II-B Methods Review for Specific Tasks
@@
OOD detection: As QA systems have evolved significantly in recent years, there is now a higher expectation for systems not only to answer known questions correctly but also to identify and reject unknown inputs, which constitutes the OOD detection task. Early OOD detection methods relied on the availability of numerous OOD data for binary classification [24]. However, the scope and variety of OOD data are theoretically infinite, making collection a burdensome task. Subsequently, more approaches have focused on training networks using only IND data to generate heterogeneous representations [25, 26]. Various downstream algorithms, such as local outlier factor (LOF) [27], Gaussian discriminant analysis (GDA) [28], energy score [29, 30], and Mahalanobis distance [31, 32], have been applied to distinguish between IND and OOD text features. The latest advances in OOD detection fall into two broad categories: using ConL methods to optimize the representation space [33, 34, 15], and employing generative modeling approaches like training a GPT model [35] on IND data to achieve lower likelihoods for OOD text generation [36, 37, 38].
@@
New intent discovery, NID: Initially, new intent discovery focused on identifying new categories directly from unlabeled data, with most methods applying unsupervised clustering [39, 40]. Later approaches commonly incorporated some labeled data to support the discovery of unknown intents [41, 42]. A typical workflow involves computing textual similarities between unlabeled and known data, assigning pseudo labels to unknown data based on the results, and then performing clustering. Recently, NID research has also emphasized the importance of representation learning, with many studies incorporating pre-training and contrastive learning methods to optimize feature space arrangements [43, 44, 45].
@@
Continual learning: Continual learning denotes a re-training procedure to let the network learn new knowledge in stages, and finally reach a level to handle all tasks. For example, we may let the model learn the even five handwriting numbers and then the odd fives, finally let it recognize all ten numbers. The biggest challenge of continual learning is the forgetting problem [46], that is, the model’s ability to recognize old data may deteriorate significantly after learning new data. Current studies are also focusing on solving this issue, where the solutions mainly include weight regularization [47, 48] and knowledge distillation [49, 50]. In this paper, for simplicity, we apply a very straightforward method that re-trains the model with both a few new and old intent data, which can be regarded as the “replay” policy [51, 52], to mitigate forgetting. Experiments (Section VI-E) show that this policy is effective and exhibits minimal performance degradation for old data classification.
@@
Despite the above significant progress in each task, there remains a gap in integrating these tasks into a comprehensive QA system building, which is the aim of this paper to realize using SCL.
III Problem and Task Definition
@@
As mentioned in Section I, we introduce four basic tasks to build a comprehensive and adaptive QA system (Figure 1). Here, before presenting the specific solutions, we formally define the problems and targets of each task.
@@ III-A Preliminary
Encoder: We have a language encoder ENC that encodes and transforms a discrete text $\bm{x}$ into an m-dimension continuous representation: $\bm{x}$$\mathbb{R}^{m}$. Since there are many text similarity calculations during SCL training, we adopt a simple but effective “mean-pooling” strategy to obtain the sentence representation. Specifically, the average of all word embeddings in the sentence is served as the sentence embedding:
@@ III-B T-1: User intent classification
For traditional methods like CE-fine-tuning, this is achieved by further adding a fully-connected layer to the end of $ENC$ and training them together using CE loss. In the proposed method, the encoder is pre-trained using SCL (refer to Section IV-A for details) and the parameters are fixed in the following tasks T-1/2/3. $f_{cls}$ is realized by a nearest neighbor algorithm (Section IV-B1).
@@ III-C T-2: Out-of-Domain detection
The target of this task is to determine whether an example $\bm{x}_{j}$ from $\mathcal{S}_{unlabeled}$ belongs to $\mathbb{D}_{IND}$ or $\mathbb{D}_{OOD}$. This is usually realized by comparing the detection score $sco$ given by a detection algorithm $f_{det}$ with an empirically-set threshold $\lambda$:
@@
Here, the detected $\bm{x}_{j^{\prime}}^{ood}$ are saved for T-3 (new intent discovery), while the detected $\bm{x}_{j^{\prime}}^{ind}$ will be classified normally in T-1 and also saved for the final task, T-4 (continual learning).
@@ III-D T-3: New intent discovery
Once the collected OOD texts $\bm{x}_{j^{\prime}}^{ood}$ reach a sufficient quantity, an algorithm $f_{nid}$ is used to cluster and assign pseudo labels $\hat{y}_{j^{\prime}}$ to them, as the “new IND intents” in $\mathcal{C}_{kn^{\prime}}$:
@@ IV Methodology
In this section, we introduce the methodology of SCL as the RepL method in the pre-training stage, followed by the downstream solutions to each task.
@@ IV-A1 Principle and optimizing effect of SCL
Specifically, as illustrated in Figure 2, for the local aspect, compared with CE fine-tuning, SCL applies a more straightforward yet effective target: pulling the representations of similar examples (positive examples) together and pushing dissimilar ones (negative examples) further apart. To tackle the global distribution problem, SCL further generates “view” augmentations as additional positive examples before the contrasting process, optimizing features in a boarder scale. Together, these strategies provide a better supervision signal to arrange text representations to form compact and distinct clusters, enhancing both the alignment and uniformity of feature space, and thus sharpening the distinctions between texts from various intents.
This well-organized embedding space (Figure 2-b) provides a strong foundation that benefits all downstream tasks:
@@ 2. T-2:
    The compact and distinct IND clusters highlight the differences between them and OOD texts, which do not belong to any known intents.
@@ 3. T-3:
    Although not trained directly with OOD data, the two optimization directions of SCL still drive the aggregation of semantically similar features. This creates an environment conducive to intent discovery algorithms, making it easier to identify the clustered gathered data.
@@ 4. T-4:
    The accurately discovered new intent data, pre-organized feature space, and SCL-based re-training, together ensure the high-performance of the updated system.
@@ IV-A2 Generation of text “views”
For images, view generation can be achieved by simple transformations such as rotating and cropping. However, these techniques are not easily applicable to text “view” generation, as hasty word swaps or deletions can distort the original meaning. In our previous study [15], we found that encoding a sentence twice with different drop-out masks (Figure 2-left) is an effective way to generate high-quality, slightly different but main meaning preserved text embeddings. By using this generation method, SCL gave the best OOD detection results compared to no-view-generation [33] or more complex techniques like embedding-attack [53] (refer to Tables 3-5 of [15] for detailed results). Considering these advantages, we continue to use this technique as the augmentation method in this paper.
@@ IV-A3 Formal learning procedure definition
$\bm{h}_{2i^{\prime}}=ENC_{drop1}(\bm{x}_{i}),\quad\bm{h}_{2i^{\prime}+1}=ENC_{drop2}(\bm{x}_{i})$ (7)
@@
IV-B Stage 2: Solutions for Downstream Tasks
With a well-organized embedding space achieved through SCL, we can complete all the tasks with the least efforts.
@@ IV-B1 Solution for T-1 intent classification
We apply a Mahalanobis distance (MDist) - nearest neighbor (NN)-based method for IND intent classification, which does not require any additional training. We use MDist because, compared to Euclidean distance, MDist considers the distribution of each feature component, providing a more accurate measurement of text similarity. Additionally, to further accelerate the classification but avoid the instability of NN, we propose calculating MDist between the current test sample $\bm{x}_{test}$ and only the centroids $\bar{\bm{x}}_{c}$ of each intent $c$. Finally, $y_{test}$ is determined as the label of the nearest centroid. The whole process is divided into three steps:
@@
The class centroid is the average data point of all text embeddings belonging to this intent:
@@
where $num(c)$ is data quantity in category c, and $\bm{h}_{c,i}$ is the text representation of its i-th example.
@@
$\small S=\frac{1}{\left|C\right|}\sum_{c\in{\mathcal{C}}}(\bm{H}_{c}-\overline{\bm{H}}_{c}){(\bm{H}_{c}-\overline{\bm{H}}_{c})}^{T}$ (10)
@@
$\small y_{\text{test}}=\arg\min_{c\in\mathcal{C}}MDist(\bm{h}_{\text{test}},\bar{\bm{h}}_{c})$ (12)

### v2

@@ I Introduction
Figure 1: The workflow of the proposed comprehensive and adaptive QA system. First, an initial model is trained using current known data (Pre). Then, during daily usage, an inquiry will first be judged to be an answerable input or not (T-2): if so, proceed to the intent detection (T-1) and retrieve the answer; if not, the unknown inquiry will be temporally saved and be clustered later (T-3). Finally, the initial model is re-trained with newly discovered data to obtain the classification ability to new intents (T-4).
@@ 1. $\bullet$
    Externally, or at the system level: Current QA systems typically adopt a “training-as-complete” policy, meaning that once the model has finished training, its knowledge is not updated. This approach leaves the system vulnerable to unseen, unknown inputs (as out-of-domain, OOD inputs [11]). To maintain the robustness and security of the model, there should be mechanisms to handle these inputs. Furthermore, even when new data are collected to upgrade the system, the common practice is to re-train a new model from scratch, leading to a waste of labor and resources.
@@ 4. T-4
Figure 2: Diagrams of SCL’s representation optimization procedure, and solution for each task. Left: the training and optimization effect by SCL, which gathers same-intent text features together and pushes different features further apart, providing an underlying embedding space for following tasks. Right: the classification/detection processes of T-1 to T-4 after SCL. (Together with Figure 1 for better understanding)
More specifically, the advantage of SCL over CE fine-tuning lies in its ability to enhance both learning efficiency and feature separation. As shown in Figure 2-b, SCL employs a simple but clear optimization target: pulling same-label text features together and pushing different ones further apart, resulting in an intra-class compact and inter-class scattered, well-organized embedding space. This clear separation between clusters facilitates easier in-domain (IND) intent classification and IND/OOD separation and detection (Figure 2, T-1/2). Additionally, this structure also supports the aggregation of currently unknown but semantically similar features, creating a conducive environment for new intent discovery and continual learning (Figure 2, T-3/4). Consequently, even with the simplest algorithms in downstream tuning, all task objectives can be achieved effortlessly and efficiently.
@@
    Our experiments demonstrate that, despite its conciseness and simplicity, the proposed solution significantly improves global feature optimization efficiency and outperforms state-of-the-art methods across all tasks, complementing the deficiencies of current QA systems. [^1] Code is available at https://github.com/bowang-rw-02/scl-one-for-four
@@ II Related Works
II-A Representation learning and intent classification model building in QA system
As introduced in Section I, the efficiency of a QA system heavily depends on the performance of the intent classification module, where the representation learning (RepL) process plays a key role in it. Here, a good RepL method can capture the intrinsic patterns and structures of the data, and thus providing a solid foundation for downstream classification and other tasks. Early RepL research focused on manually extracting meaningful features from raw data, with preliminary methods including TF-IDF [4] and BoW [5]. Subsequently, the development of word embeddings extracted via neural networks automated the feature extraction process [6, 7]. More recently, with the advent of large-scale language encoders like BERT [8] that are pre-trained on extensive corpora, obtaining high-quality text embeddings has become much easier. The common classification model building approach is to use BERT as the base encoder, with CE fine-tuning as the RepL method, which is widely used in various recent QA systems [16, 10, 1, 17].
Drawback of BERT+CE mode
However, this BERT+CE paradigm is proven to be suboptimal for feature optimization these years. Firstly, from a global perspective, [18, 12] reported that the overall text embedding distribution generated by BERT is anisotropic. This means that most embeddings are spatially distributed in an elongated cone-like space rather than uniformly in a sphere, leading to high similarity between any two sentence representations and thus posing challenges for downstream tasks. Secondly, from a local perspective, CE fine-tuning usually results in a “loose” feature arrangement [13, 14], where the boundaries between different intent clusters are barely distinguishable. This not only reduces the performance of known, IND intent classification, but also makes the model prone to confusing IND inputs with future unseen but similar OOD inputs.
@@ II-A Representation learning and intent classification model building in QA system
To overcome above limitations and further improve learning efficiency, the contrastive learning (ConL), has been actively studied recently. Earliest ConL can be traced back to supervised metric learning (MeL) [19] used in image classification tasks. Different with the vague and loose optimization goal of CE fine-tuning mode, MeL applies a simpler but straightforward target of “gathering similar examples together and pushing dissimilar ones further apart”. This helps generate compact feature clusters efficiently, and established the foundation of ConL. Subsequently, unsupervised ConL (UConL) [20, 21] was proposed, sharing the same strategy but aiming to improve the overall representation uniformity. Finally, SCL [14] incorporates the advantages of both MeL and UConL, thus having the potential to address both global and local feature optimization issues.
@@
Despite these advantages, currently ConL is still mainly examined in computer vision (CV) field. In natural language processing (NLP) field, there are only limited applications, primarily in natural language inference (NLI) tasks using UConL methods [22, 23]. A boarder exploration of SCL’s learning effect across various tasks, including its application in QA systems, remains unaddressed, which is exactly our main research target this time.
II-B Past work on individual tasks
To provide a boarder perspective, we briefly review previous methods for addressing the three newly introduced tasks.
@@ II-B Past work on individual tasks
T-2 Out-of-Domain, OOD detection: Although QA systems have evolved significantly in recent years, they can only work correctly on the known inquires and lack the ability to recognize and reject unknown inputs, which raises the issue of OOD detection [11]. To solve this problem, early methods built a binary IND/OOD detection system using numerous OOD data [24]. However, the scope and variety of OOD data can be infinite, making this method impractical. Subsequently, more approaches have focused on training networks using only IND data to generate heterogeneous representations [25, 26], then applied various downstream algorithms, such as local outlier factor (LOF) [27], Gaussian discriminant analysis (GDA) [28], energy score [29, 30], and Mahalanobis distance [31, 32], to distinguish IND and OOD text features. Latest advances fall into two broad categories: using ConL methods to optimize the representation space [33, 34, 15], or employing generative modeling approaches like training a GPT model [35] on IND data to achieve lower likelihoods for OOD text generation [36, 37, 38].
@@
T-3 New intent discovery, NID: NID is a classic and substantially studied task in NLP. Early studies focused on identifying new categories directly from unlabeled data, with most methods applying unsupervised clustering [39, 40]. The later approaches commonly incorporated some labeled data to support the discovery of unknown intents [41, 42]. A typical workflow involves computing textual similarities between unlabeled and known data, assigning pseudo labels to unknown data based on the results, and then performing clustering. Recently, NID research has also emphasized the importance of RepL, with many studies incorporating pre-training and contrastive learning methods to optimize feature space arrangements [43, 44, 45].
@@
T-4 Continual learning: Continual learning generally denotes a re-training procedure to let the network learn new knowledge in stages, and finally reach a level to handle all tasks. For example, we may let a model learn the even five handwriting numbers and then the odd fives, finally let it recognize all ten numbers. The biggest challenge of continual learning is catastrophic forgetting [46], where the model’s performance on earlier tasks deteriorates as it learns new tasks. Current studies have also been focusing on solving this issue, where the solutions mainly include: a) weight regularization [47, 48], which imposes some constraints on the model’s parameters to preserve values of model weights associated with knowledge from old data; b) knowledge distillation [49, 50], which builds a smaller model to mimic the past model output as the hint for old tasks during handling new tasks; c) replay [51, 52], which involves storing a subset of old data and replaying it during training on new tasks to reinforce the old knowledge.
@@
Although above methods have made much progress in their respective tasks, there still remains a gap in systematically integrating these tasks into a comprehensive QA system building, which is the aim of this paper.
III Task and workflow definition
@@
As mentioned in Section I, we introduce four basic tasks to build a comprehensive and adaptive QA system (Figure 1). Before presenting the specific solutions in the next section, we formally define these tasks and also simply introduce the workflow for better comprehension.
@@ III-A Preliminary
Encoder: We have a language encoder ENC that encodes and transforms a discrete text $\bm{x}$ into an m-dimension continuous representation $\bm{h}$: $\bm{x}$$\mathbb{R}^{m}$. Since there are many text similarity calculations during SCL training, we adopt a simple but effective “mean-pooling” strategy to obtain the sentence representation. Specifically, the average of all word embeddings in the sentence is served as the sentence embedding:
@@ III-B T-1: User intent classification
For traditional methods like CE-fine-tuning, this is achieved by further adding a fully-connected layer to the end of ENC and training them together using CE loss. In the proposed method, ENC is pre-trained using SCL (refer to Section IV-A for details) and the parameters are fixed during T-1 and the following T-2 and T-3. The $f_{cls}$ is realized by a nearest neighbor algorithm (Section IV-B1).
@@ III-C T-2: Out-of-Domain detection
The target of this task is to determine whether an example $\bm{x}_{j}$ from $\mathcal{S}_{unlabeled}$ belongs to $\mathbb{D}_{IND}$ or $\mathbb{D}_{OOD}$. This is usually realized by comparing the detection score $sco$ of $\bm{x}_{j}$ with an empirically-set threshold $\lambda$, where $sco$ is given by a detection algorithm $f_{det}$ analyzed on the text embedding:
@@
In the actual workflow (as shown in Figure 1), the detected $\bm{x}_{j^{\prime}}^{ood}$ samples are saved for T-3 (new intent discovery), while the detected $\bm{x}_{j^{\prime}}^{ind}$ examples will be classified normally in T-1 and also saved for the final task, T-4 (continual learning).
@@ III-D T-3: New intent discovery
Once the collected OOD texts $\bm{x}_{j^{\prime}}^{ood}$ from T-2 reach a sufficient quantity, an algorithm $f_{nid}$ can be used to cluster and assign pseudo labels $\hat{y}_{j^{\prime}}$ to these examples, as the “new IND intent” labels stored in set $\mathcal{C}_{kn^{\prime}}$:
@@ IV Methodology
In this section, we introduce the methodology of SCL as the RepL method in the pre-training stage, followed by the downstream methods to solve to each task.
@@ IV-A1 Principle and optimizing effect of SCL
Specifically, as illustrated in Figure 2-b, for the local aspect, compared with CE fine-tuning, SCL applies a more straightforward yet effective target: pulling the representations of similar examples (as positive examples) together and pushing dissimilar ones (as negative examples) further apart. To tackle the global distribution problem, SCL further generates “view” augmentations as additional positive examples before the contrasting process (Figure 2-Pre-1), optimizing features in a boarder scale. Together, these strategies provide a better supervision signal to arrange text representations to form compact and distinct clusters, enhancing both the alignment and uniformity of the feature space, and thus sharpening the distinctions between texts from various intents.
This well-organized embedding space provides a strong foundation that benefits all downstream tasks (as Figure 2-right part shows):
@@ 2. T-2:
    The compact and distinct IND clusters highlight the differences between them and OOD texts that do not belong to any known intents.
@@ 3. T-3:
    Although not trained directly with OOD data, the two optimization directions of SCL still drive the aggregation of semantically similar features. This creates an environment conducive to intent discovery algorithms, making it easier to identify the pre-gathered data.
@@ 4. T-4:
    The accurately discovered new intent data, pre-organized feature space, and SCL-based re-training, together ensure the continuous high-performance of the updated system.
@@ IV-A2 Generation of text “views”
For images, view generation can be achieved by simple transformations such as rotating and cropping. However, these techniques are not easily applicable to text “view” generation, as hasty word swaps or deletions can distort the original meaning. In our previous study [15], we found that encoding a sentence twice with different drop-out masks (Figure 2-Pre-1) is an effective way to generate high-quality, slightly different but main meaning preserved text embeddings. By using this generation method, SCL gave the best OOD detection results compared to no-view-generation [33] or more complex techniques like embedding-attack [53] (refer to Tables 3-5 in [15] for detailed results). Considering these advantages, we continue to use this technique as the augmentation method in this paper.
@@ IV-A3 Formal learning procedure definition
$\bm{h}_{2i^{\prime}-1}=ENC_{drop1}(\bm{x}_{i}),\quad\bm{h}_{2i^{\prime}}=ENC_{drop2}(\bm{x}_{i})$ (7)
@@
IV-B Stage 2: Solution for each task after SCL
With a well-organized embedding space achieved through SCL, now all four downstream tasks can be solved using simple methods with the least efforts.
@@ IV-B1 Solution for T-1 intent classification
For T-1, as shown in Figure 2-T-1, we apply a Mahalanobis distance (MDist) - nearest neighbor (NN)-based method to classify the text representations encoded by SCL-trained ENC, which does not require any additional training. We choose MDist because, compared to Euclidean distance, MDist considers the distribution of each feature component, providing a more accurate measurement of text similarity. Additionally, to further accelerate the classification speed but avoid the instability of NN, we propose calculating MDist between the current test sample $\bm{x}_{test}$ and only the centroids $\bar{\bm{x}}_{c}$ of each intent $c$. Finally, the label of one test example $y_{test}$ is determined as the label of the nearest centroid. The whole process can be described into the following three steps:
@@
The class centroid of category c is the average data point of all text embeddings belonging to this intent:
$\bm{h}_{c,i}=ENC(\bm{x}_{c,i})$ (9)
@@
where $num(c)$ is data quantity of category c, and $\bm{h}_{c,i}$ is the text representation of its i-th example $\bm{x}_{c,i}$.
@@
$\small S=\frac{1}{\left|{\mathcal{C}}_{kn}\right|}\sum_{c\in{\mathcal{C}_{kn}}}(\bm{H}_{c}-\overline{\bm{H}}_{c}){(\bm{H}_{c}-\overline{\bm{H}}_{c})}^{T}$ (11)
@@
$\small y_{\text{test}}=\arg\min_{c\in\mathcal{C}_{kn}}MDist(\bm{h}_{\text{test}},\bar{\bm{h}}_{c})$ (13)